Consider a vertical spring with spring constant 23
Answer #1


using law of conservation of energy


energy at the 8.815 cm = energy at the maximum distance from equilibrium


0.5*k*x^2 + 0.5*m*v^2 = 0.5*k*A^2


(0.5*23.45(8.815*10^-2)^2) + (0.5*1.007*2.58^2*10^-2) = 0.5*23.45*A^2

A = 0.10309 m = 1.03*10^-1 m


so correct option is H) 1.03*10^-1 m